OpenStudy (anonymous):

Sorry to bother.Tatiana is designing a label for her band's new CD. The label is circular in shape with a hole in the middle. If the label has a diameter of 4.5 inches, and the center hole has an area of 1.8 square inches, what is the area she has for her design? Use π = 3.14. Round the answer to the nearest hundredth.

OpenStudy (anonymous):

\[3.14(4.5)^2-3.14(1.8)^2\]

OpenStudy (anonymous):

oh okay I was doing the problem wrong thank you so very much

OpenStudy (phi):

The radius is 4.5/2 = 2.25 inches the area of the CD is \[ A= \pi r^2 \] A= 3.14*2.25*2.25= 15.896 sq in the inner hole has area 1.8 sq inches, which we subtract away: A= 15.896-1.8= 14.096 round to the nearest hundred A= 14.10